OpenStudy (anonymous):

If you only had a derivative of a function, and you had to work out what the plot of the original curve was, and maybe the derivative too, how would you approach that problem?

OpenStudy (michele_laino):

what is: E^2x

OpenStudy (anonymous):

I just made it up.. how about E^x

OpenStudy (anonymous):

\[f'[x]=(x+2)/E^x\]

OpenStudy (michele_laino):

we can integrate yor first derivative, nevertheless I don't know how, since I don't know the meaning of E^x

OpenStudy (anonymous):

is that not E to the power of x?

OpenStudy (michele_laino):

yes! Is E the base of natural logarithms

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So I would need to integrate.. I haven't learned that yet.. I guess that's coming next..

OpenStudy (michele_laino):

ok! then we have to compute this integral: \[\Large \int {\left( {x + 2} \right){e^{ - x}}} dx\]

OpenStudy (michele_laino):

after a simple computation, we get: \[\Large \int {\left( {x + 2} \right){e^{ - x}}} dx = - {e^{ - x}}\left( {x + 3} \right) + C\] where C is the usually arbitrary real constant

OpenStudy (michele_laino):

in other words you have a family of functions, which differ each other by an additive constant C